# Examining the Source of Defects from a Mechanical Perspective for 3D Anomaly Detection

### v1

@@ 6 Conclusion
This paper presents MC4AD, a novel 3D anomaly detection method that investigates the source of anomalies from a mechanical perspective. By considering anomalies as the result of the action of a damaging force, we repaired the anomalies utilizing a corrective force. We generate diverse anomalies by DA-Gen and obtain capture of potentially differentiable damage forces using the proposed CFP-Net to predict the corresponding corrective forces. And we guarantee the correct generation of the corrective force through new combined loss functions. Moreover, a dataset with intraclass variance named Anomaly-IntraVariance and a hierarchical quality control strategy were proposed to further simulate a real assembly line environment. Nine advanced anomaly detection and segmentation performance metrics are demonstrated in extensive comparative experiments on the proposed and existing five datasets across 90 categories. Limitation. MC4AD lacks explicit physical constraints to derive damage forces for future work; we discuss more in Appendix H.
@@ A.2 How We Select Baseline
We select models according to their performance, and some lower-performing models are excluded from the selection. Moreover, some excellent work based on large language models has excelled in 3D anomaly detection. Considering the large amount of data they use that is difficult to access in the real industry, we compare models trained with limited data, such as migration training using PointMAE, or on local datasets, such as those trained on Real3D-AD. We state here some of the models that were not selected for comparison for this reason.
@@
The above work is excellent, and we collected it in Arxiv or the official conference website. We discuss them in related work, and due to the task setup, we do not impose comparison experiments.
@@ A.3 Implementation of Baselines
IMRNet eliminates potential anomalies in the point cloud by expanding PointMAE to circularly mask the reconstruction. Since its code is not open-sourced, we obtained its official paper performance through the following official repository: https://github.com/Chopper-233/Anomaly-ShapeNet.
R3D-AD eliminates potential anomalies in the point cloud by diffusion modelling. Since it is not open source, we get the official performance at the link: https://arxiv.org/abs/2407.10862v1.
ISMP obtains information about the interior of the point cloud through pseudo-modal projections as complementary information to obtain better anomaly detection, and we access the official implementation through the following link: https://github.com/M-3LAB/Look-Inside-for-More.
@@ A.4 Details of the HQC Generalisation to Memory Bank Methods
The core of HQC is an initial screening with a simplified version that uses its faster but slightly reduced detection performance to detect a fraction of significantly normal samples, leaving the other fraction of potentially abnormal samples for further detection using the original model. When the model is properly initialised, it speeds up detection without loss of accuracy.
@@ A.6 Details of Generalizability Experiments
Following the modified PointNet of Eq. 13, the DGCNN inserts dynamic graph aggregation after each convolution. Moreover, PointNet++ feature extractor for multi-scale point clouds using PointNet.
@@ Appendix B More Analysis about Motivation
In this paper, we view most of the anomalies that arise in industrial manufacturing as the combined result of internal and external damage forces. Moreover, we further elaborate on the theoretical implications of the individual formulas and relate them to 3D anomaly detection.
@@ B.1 Analysis of Potential Continuous Damage Force and Correction Force
Continuous forces are usually expressed as integrals of force densities over a spatial distribution. Consider that mass $M$ can be written as a product of density $\rho$ and volume $V$ phy . In 3D-AD, the basic unit for detecting samples is the point cloud. Therefore, we ignore the mass density of the object, and each point can be ideally regarded as a unit density. Furthermore, as the point cloud is composed in the form of discrete points, the volume $V$ can be approximated as a continuous surface $S$. The force representation in the 3D point cloud can be written:
@@
Eq. 15 represents the combined action of complex resultant forces on the surface $S$. However, it is straightforward to make the model learn the complex potential resultant forces. Therefore, we further decompose the resultant forces to obtain the external force $F_{E}$ and internal force $F_{I}$, which can be represented as follows:
@@ B.2 Analysis of Different Forces to Normal Points and Anomaly Points
Eq. 18 can be derived from Newton’s First Law of Motion. if the normal structure is undamaged, which means that the point is subjected to zero damaging force, and we do not need to apply a corrective force to rectify it. On the contrary, if a non-zero damage force is suffered, this means that the point is anomalous and, according to the definition in Definition A4, we need to apply an opposite corrective force to repair it.
@@
Since our model learns potential damage forces from the point cloud and generates corrective forces to counteract the effects of the damage forces through an inverse process. This process behaves as an asymmetric inverse process. The model is required to solve the following challenges:
@@
For Challenge A2. As explicitly formulated in Eq. 8 through complementary constraints, our model employs the MC-Skip mechanism to adjust feature weights across different hierarchies dynamically. This design guides the model to learn complementary feature representations, which are further regularized by a symmetric loss function as mentioned in Eq. 10.
@@
Since the whole process is microscopic, this leads to our model being more suitable for learning potentially microscopic variables.
@@ D.1 Data Overview
We constructed an anomaly detection dataset with intra-class variance as shown in Figure 5. Group 1 contains eight classes, where each class has two subclasses. Group 2 also contains eight classes, where each class has four subclasses. Moreover, to more closely mimic 3D anomaly detection in a real industrial context, Group 2 has a smaller number of points and has been subjected to light noise. This is attributed to the tens of minutes required to scan a high-resolution sample such as Real3D-AD, which may pose challenges in the real industry. Thus, we provide a more comprehensive test benchmark for 3D anomaly detection in the real industry.
@@ D.2 Anomaly Data Synthesis
The data contains more than 1,000 samples in 16 categories, four samples in the training set, and dozens of samples in the test set, and the statistics for Group 1 and Group 2 are shown in Tables 6 and 7, respectively. Sample point clouds range from 7,000 to 25,000 points, meeting the scanning accuracy of the actual industry.
@@ E.2 Parameter Sensitivity Analyses to DA-Gen
DA-Gen contains three key parameters: $\lambda$, $\sigma$, and $\gamma$. We set a reference value based on empirical evidence and show the results of the parameter sensitivity analyses in Tables 8(b), 8(c), and 8(d), respectively. A smaller $\lambda$ indicates that the direction in which the anomaly occurs is shifted towards the random direction to a greater extent. When $\lambda$ was set to a smaller [0.900,1] and larger [0.975, 1] resulted in an average decrease of 1.62% and 2.25% for O-AUROC and P-AUROC, respectively. We select a moderate value $\lambda=$[0.950,1] to generate the true anomaly. The stretching parameter $\sigma$ is used to control the degree of narrowing of the anomalous shape. We chose $\sigma=$[0,0.08] to obtain the best performance for the anomaly detection metrics of 78.6% and 83.7% compared to the performance degradation of 2.25% and 2.95% caused by the narrower $\sigma=$[0,0.02] and the wider $\sigma=$[0,0.16]. $\gamma$ represents the maximum displacement magnitude of the center of the anomaly. The $\gamma$ set to a larger [0.09,0.15] or smaller [0,0.06] resulted in an average decrease in the anomaly detection metrics of 5.5% and 6.05%. We choose $\gamma$=[0.06,0.12] for optimal anomaly creation. We simply visualise the generated point cloud in Appendix F.3.
@@ Appendix H Limitations and Feature Work
We discuss further limitations in this section and point the direction of future work. Our framework lacks explicit physical constraints. Current approaches in the robotics area incorporating physical engines barcellona2025dream (e.g., LLM-based priors) face challenges in 3D anomaly detection, primarily due to the inherent structural ambiguity of 3D point clouds and the non-physical randomization in synthetic dataset generation. Notably, real-world benchmarks we use like Real3D-AD liu2023real3d and MvTec3D-AD Bergmann_2022 maintain authentic physical constraints. Future dataset development should integrate comprehensive physical descriptors (e.g., RGB textures, illumination conditions, and industrial contexts) to enable explicit physical regularization in anomaly detection models.
@@
Table 18: O-AUROC performance of different methods on Real3D-AD across 12 categories, where best and second-place results are highlighted in red and blue, respectively.
O-AUROC

### v2

@@ 6 Conclusion
This paper presents MC4AD, a novel 3D anomaly detection method that investigates the source of anomalies from a mechanical perspective. By considering anomalies as the result of the action of a damaging force, we repaired the anomalies utilizing a corrective force. We generate diverse anomalies by DA-Gen and obtain capture of potentially differentiable damage forces using the proposed CFP-Net to predict the corresponding corrective forces. We also guarantee the correct generation of the corrective force through new combined loss functions. Moreover, a dataset with intraclass variance named Anomaly-IntraVariance and a hierarchical quality control strategy were proposed to simulate a real assembly line environment further. Extensive comparative experiments on the proposed and existing five datasets across 90 categories demonstrate nine advanced anomaly detection and segmentation performance metrics. Limitation. MC4AD lacks explicit physical constraints to derive damage forces for future work; we discuss more in Appendix H.
@@ A.2 How We Select Baseline
We select models according to their performance, and some lower-performing models are excluded from the selection. Moreover, some excellent work based on large language models has excelled in 3D anomaly detection. Considering the large amount of data they use that is difficult to access in the real industry, we compare models trained with limited data, such as migration training using PointMAE, or local datasets, such as those taught on Real3D-AD. For this reason, we state some models that were not selected for comparison.
@@
The above work is excellent, and we collected it in Arxiv or the official conference website. We discuss them in related work, and we do not impose comparison experiments due to the task setup.
@@ A.3 Implementation of Baselines
IMRNet eliminates potential anomalies in the point cloud by expanding PointMAE to mask the reconstruction circularly. Since its code is not open-sourced, we obtained its official paper performance through the following official repository: https://github.com/Chopper-233/Anomaly-ShapeNet.
R3D-AD eliminates potential anomalies from a point cloud by a diffusion model, and we obtained its official paper performance through the following official repository: https://github.com/zhouzheyuan/r3d-ad.
ISMP obtains information about the interior of the point cloud through pseudo-modal projections as complementary information to get better anomaly detection, and we access the official implementation through the following link: https://github.com/M-3LAB/Look-Inside-for-More.
@@ A.4 Details of the HQC Generalisation to Memory Bank Methods
The core of HQC is an initial screening with a simplified version that uses its faster but slightly reduced detection performance to detect a fraction of significantly normal samples, leaving the other fraction of potentially abnormal samples for further detection using the original model. When properly initialized, the model speeds up detection without losing accuracy.
@@ A.6 Details of Generalizability Experiments
Following the modified PointNet of Eq. 13, the DGCNN inserts dynamic graph aggregation after each convolution. Moreover, the PointNet++ feature extractor for multi-scale point clouds using PointNet.
@@ Appendix B More Analysis about Motivation
In this paper, we view most anomalies in industrial manufacturing as the combined result of internal and external damage forces. Moreover, we elaborate further on the individual formulas’ theoretical implications and relate them to 3D anomaly detection.
@@ B.1 Analysis of Potential Continuous Damage Force and Correction Force
Continuous forces are usually expressed as integrals of force densities over a spatial distribution. Mass $M$ can be written as a product of density $\rho$ and volume $V$ phy . In 3D-AD, the basic unit for detecting samples is the point cloud. Therefore, we ignore the object’s mass density, and each point can be ideally regarded as a unit of density. Furthermore, as the point cloud is composed of discrete points, the volume $V$ can be approximated as a continuous surface $S$. The force representation in the 3D point cloud can be written:
@@
Eq. 15 represents the combined action of complex resultant forces on the surface $S$. However, making the model learn the complex potential resultant forces is straightforward. Therefore, we further decompose the resultant forces to obtain the external force $F_{E}$ and internal force $F_{I}$, which can be represented as follows:
@@ B.2 Analysis of Different Forces to Normal Points and Anomaly Points
Eq. 18 can be derived from Newton’s First Law of Motion. if the normal structure is undamaged, the point is subjected to zero damaging force, and we do not need to apply a corrective force to rectify it. On the contrary, if a non-zero damage force is suffered, the point is anomalous and, according to the definition in Definition A4, we need to apply an opposite corrective force to repair it.
@@
Since our model learns potential damage forces from the point cloud and generates corrective forces to counteract the effects of the damage forces through an inverse process, this process behaves as an asymmetric inverse process. The model is required to solve the following challenges:
@@
For Challenge A2. As explicitly formulated in Eq. 8 through complementary constraints, our model employs the MC-Skip mechanism to dynamically adjust feature weights across different hierarchies. This design guides the model to learn complementary feature representations, which are further regularized by a symmetric loss function as mentioned in Eq. 10.
@@
Since the whole process is microscopic, our model is more suitable for learning potentially microscopic variables.
@@ D.1 Data Overview
We constructed an anomaly detection dataset with intra-class variance as shown in Figure 5. Group 1 contains eight classes, each class having two subclasses. Group 2 also includes eight classes, where each class has four subclasses. Moreover, to more closely mimic 3D anomaly detection in a real industrial context, Group 2 has fewer points and has been subjected to light noise. This is attributed to the tens of minutes required to scan a high-resolution sample such as Real3D-AD, which may pose challenges in the real industry. Thus, we provide a more comprehensive test benchmark for 3D anomaly detection in the real sector.
@@ D.2 Anomaly Data Synthesis
The data contains more than 1,000 samples in 16 categories, four samples in the training set, and dozens of samples in the test set, and the statistics for Group 1 and Group 2 are shown in Tables 7 and 8, respectively. Sample point clouds range from 7,000 to 25,000 points, meeting the scanning accuracy of the actual industry.
@@ E.2 Parameter Sensitivity Analyses to DA-Gen
DA-Gen contains three key parameters: $\lambda$, $\sigma$, and $\gamma$. We set a reference value based on empirical evidence and show the results of the parameter sensitivity analyses in Tables 9(b), 9(c), and 9(d), respectively. A smaller $\lambda$ indicates that the direction in which the anomaly occurs is shifted towards the random direction to a greater extent. When $\lambda$ was set to a smaller [0.900,1] and larger [0.975, 1], it resulted in an average decrease of 1.62% and 2.25% for O-AUROC and P-AUROC, respectively. We select a moderate value $\lambda=$[0.950,1] to generate the true anomaly. The stretching parameter $\sigma$ controls the degree of narrowing of the anomalous shape. We chose $\sigma=$[0,0.08] to obtain the best performance for the anomaly detection metrics of 78.6% and 83.7% compared to the performance degradation of 2.25% and 2.95% caused by the narrower $\sigma=$[0,0.02] and the wider $\sigma=$[0,0.16]. $\gamma$ represents the maximum displacement magnitude of the center of the anomaly. The $\gamma$ set to a larger [0.09,0.15] or smaller [0,0.06] resulted in an average decrease in the anomaly detection metrics of 5.5% and 6.05%. We choose $\gamma$=[0.06,0.12] for optimal anomaly creation. We simply visualise the generated point cloud in Appendix F.3.
@@ Appendix H Limitations and Feature Work
We discuss further limitations in this section and point the direction of future work. Our framework lacks explicit physical constraints. Current approaches in the robotics area, incorporating physical engines barcellona2025dream (e.g., LLM-based priors) face challenges in 3D anomaly detection, primarily due to the inherent structural ambiguity of 3D point clouds and the non-physical randomization in synthetic dataset generation. Notably, real-world benchmarks we use, like Real3D-AD liu2023real3d and MvTec3D-AD Bergmann_2022 maintain authentic physical constraints. Future dataset development should integrate comprehensive physical descriptors (e.g., RGB textures, illumination conditions, and industrial contexts) to enable explicit physical regularization in anomaly detection models.